OpenStudy (ammarah):

Write an equation of the ellipse with a vertex at (0,-10), a co-vertex at (6,0), and center (0,0) @ganeshie8

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

its a long thin ellipse

ganeshie8 (ganeshie8):

a = 10 b = 6

ganeshie8 (ganeshie8):

write out the equation

ganeshie8 (ganeshie8):

\(\large \frac{x^2}{b^2} + \frac{y^2}{a^2} = 1\) \(\large \frac{x^2}{6^2} + \frac{y^2}{10^2} = 1\)

OpenStudy (ammarah):

is that it?

OpenStudy (ammarah):

do i have to write x^/36+y^/100=1? or can i just leave it like that^